OpenStudy (anonymous):

log₅(4x+1)-log₅(x-5)=2 x=_____?

OpenStudy (anonymous):

in this similar problem, x=6 log(2)(x+2)+log(2)(x-4)=4

OpenStudy (anonymous):

When you have the same base (in this case it is 5) you can just cross out the logs. Then solve (4x+1)(x-5)=2

OpenStudy (anonymous):

okay, then I get 6/5... that doesn't look right?

OpenStudy (anonymous):

oh, it would be (4x+1)(X-5)=2^5 I believe sorry been a while since I have done logs

OpenStudy (anonymous):

Or it's =5^2 :p

OpenStudy (anonymous):

oh, thanks!

OpenStudy (anonymous):

np, look up Log rules on google, there are a bunch and they saved me in trig :P

OpenStudy (anonymous):

eek, the answer was 6 again!